Dclark28:

In one U.S.​ city, the taxi cost is 4.00 plus 0.70 per mile. If you are traveling from the​ airport, there is an additional charge of 3.00 for tolls. How far can you travel from the airport by taxi for 31.50 ​$​?

jhonyy9:

do you have any idea about this ? so like a first step you need to pay 4,00 plus 0,70 per mile but there is again an additional charge of 3,00 for tolls you have total 31,50$ how far can you travel by taxi from the airport for this 31,50$ ? how you think what may be the first step ?

surjithayer:

4+0.70x+3=31.50 0.70x=31.50-7=24.50 70x=2450 x=?